OpenStudy (anonymous):

if incoming water is 12 C and you need to raise it to 42 C, what is the maximum amount of water we can heat with a power source delivering 200 kW/hour?

OpenStudy (jamesj):

I just edited the question to make it clearer. A question however: a power source is rated in Watts. So is it the power source is 200 kW. Or is it we have 200 kW.hours of energy?

OpenStudy (anonymous):

it should be energy so as to use that to convert it to heat energy......

OpenStudy (jamesj):

In other words, 200 kW/hour makes no sense. It has to be either 200 kW or 200kW.hours. @ciaran: which is it?

OpenStudy (jamesj):

We can conceive the question of "what is the amount of water we can raise from 12 C to 42 C with 200 kW.hours of energy?" or "What is the amount of water per hour we can raise from 12 C to 42 C with 200 kW of power?" Which of these two is it?

OpenStudy (jamesj):

@ciaran, please clarify the question.

OpenStudy (anonymous):

Let the maximum amount of water that can be heated per hour is m. \[Q=m \times s \times \Delta T\]...(i) \[\Delta Q= 200 \times 1\]....(ii) from (i) and (ii) find m...

OpenStudy (anonymous):

in (i) replace Q by \[\Delta Q\]... that's mistake. and compare (i) and (ii)

OpenStudy (anonymous):

No. @Taufique please check your units. @JamesJ brings about valid points here. First, note that\[Q {[\rm J]} = m {\rm [kg]} \cdot c_p \left[ \rm kJ \over kg \cdot K \right] \cdot \Delta T {\rm [K]}\]where a Joule is of units of energy. Now, if the power plant is rated in terms of kW (one of the two options JamesJ presents)\[\dot Q {\rm \left[W ~ or ~ {J \over s}\right ]} = \dot m \left [ \rm kg \over s \right] \cdot c_p \left [\rm kJ \over kg \cdot K \right] \cdot \Delta T {\rm [K]}\] Since the power plant is rated in kW/hour, we have an imaginary unit of\[\rm energy \over time^2\]I can conceive no way to manipulate the above equation to match these units. In JamesJ's second case (kW*hour), we have units of Joules.

OpenStudy (anonymous):

@ciaran If you desire an answer, you need to clarify the units on the power source.

OpenStudy (anonymous):

@eashmore ,here time is not given ,hence we can find only the amount of water in kg/time unit.